# Visual Structures Helps Visual Reasoning: Addressing the Binding Problem in VLMs

### v2

@@ 5 Discussion
This research investigated the binding problem in VLMs, a key factor limiting their performance on structured visual reasoning tasks. We introduced a method combining simple visual scaffolding, using horizontal lines, with a targeted textual prompt to encourage sequential, spatially-aware processing. Our experiments demonstrate that this intervention leads to substantial and consistent performance improvements across visual search, counting, scene description, and spatial relationship understanding. These results support the hypothesis that explicit low-level visual structure improves feature-object binding.
@@
A central finding is the critical role of visual structures. While textual prompts alone proved insufficient and sometimes detrimental, the addition of visual lines significantly improved reasoning accuracy. This suggests that current VLMs, despite their advanced linguistic capabilities, can benefit greatly from external visual cues that facilitate a more systematic parsing of complex scenes. The visual scaffolding appears to help models approximate a serial attentional process, mitigating the interference and illusory conjunctions when processing multiple objects and their features in parallel. This localized, sequential processing may also reduce representational interference by effectively managing feature entropy within sub-regions of the image, a factor highlighted in [9] as critical for robust visual reasoning. This supports the idea that addressing the binding problem may require interventions at the level of visual input processing, rather than relying solely on linguistic instruction.
@@
The proposed method offers a practical and efficient intervention, readily applicable due to its simplicity, negligible computational overhead, and single-query operation, without reliance on external tools, model fine-tuning, or architectural changes. This highlights visual input structuring as a key strategy for enhancing VLM reasoning, beyond linguistic instruction. While effective, the current static nature of the visual scaffolding represents a potential area for refinement. Future research could investigate adaptive visual scaffolding, where cues are dynamically determined by image content or query specifics. Moreover, developing VLM architectures that inherently support serial, spatially grounded attention, perhaps inspired by these findings, could offer more integrated and robust solutions to the binding problem.
@@
Last but not the least, we emphasize that improved visual reasoning in AI systems can benefit a wide range of applications, from assistive technologies and robotics to education and healthcare. However, as VLMs become more capable, their potential misuse—for instance, in surveillance, automated decision-making, or misinformation—also grows. Our approach does not require model retraining, making it widely accessible, but this also implies that such modifications could be adversarially deployed in black-box systems without transparency or oversight. To mitigate risks, we advocate for responsible use of these techniques in alignment with accountability and transparency guidelines. Moreover, future research should profoundly study the unintended biases that scaffolding might introduce in tasks other than spatial reasoning, especially when used in safety-critical applications.
@@ Appendix A Ablation study
In our ablation study, we assess multiple scene-structuring strategies and their impact on task performance using the GPT-4o [39] model. All experiments are conducted on a consistent subset of our 2D synthetic dataset [9], with varying object counts per scene. The results in Table 5 show the average performance across these scenes.
Table 5: Comparison of our method and other ablation methods on GPT-4o in 2D scenes across multiple tasks.
Method Scene Desc. (Edit Dist.) Counting (Accuracy) Visual Search (Accuracy)
Simple 1.81 16.85 0.48
Ours 1.61 31.55 0.73
Ours–NoRowNum 1.71 19.14 0.70
Column 2.00 16.68 0.76
Grid 2.43 17.40 0.72
The Simple setting represents the baseline configuration, where the model is prompted to describe the image without any additional structuring cues. As shown, this baseline yields the lowest or near the lowest accuracy in counting and visual search, indicating that unstructured prompts are insufficient for supporting systematic reasoning.
@@
In contrast, our row-based horizontal structuring, which explicitly divides the image into horizontal bands, leads to improved performance across all tasks. This design encourages the model to scan the scene in a consistent, line-by-line fashion. Furthermore, removing numerical indices from rows (Ours–NoRowNum) results in a substantial drop in accuracy, highlighting the value of explicit ordering cues in guiding sequential attention.
These results collectively underscore the importance of structured, order-aware visual annotations in enabling VLMs to reason more effectively over complex spatial arrangements.
A.1 Reasoning trace evaluation
@@
Figure 5: Comparison between our method and GPT-4o in the 2D scene description task, where models sequentially scan the image row by row to identify objects. Results are reported for scenes with 10, 15, and 20 objects using three evaluation metrics: F1 score, Jaccard similarity, and Edit distance.
Appendix B Comparison with RL-finetuned Model
We conduct a comprehensive comparison with OpenVLThinker [43], which enhances Qwen2.5-VL through an iterative training pipeline combining supervised finetuning (SFT) and reinforcement learning (RL) to strengthen multimodal reasoning. In contrast, our method requires no post-training phase and operates entirely during inference, yet achieves superior performance across most evaluated tasks. As demonstrated in Figure 6, our approach consistently outperforms both the baseline Qwen2.5-VL and OpenVLThinker in most reasoning tasks, with particularly significant gains of up to 26% on 2D-Counting task. The only exceptions where OpenVLThinker shows an advantage are in 2D-Spatial reasoning and 3D-Counting tasks, a likely consequence of its RL-based optimization for multimodal reasoning. This suggests that while training-intensive methods can excel in specific domains, they come with substantial computational costs. Our training-free paradigm, by contrast, maintains broader dominance across diverse multimodal reasoning tasks while offering superior scalability and efficiency.
Figure 6: Comparison of our method with a finetuned visual reasoning model on Qwen2.5-VL as the base model across four tasks. Each subplot shows the performance of three methods—Qwen-Simple (Qwen2.5-VL), Qwen-Ours (our method applied to Qwen2.5-VL), and OpenVLThinker (Qwen2.5-VL finetuned with RL-based reasoning)—evaluated using a task-specific metric indicated on the x-axis Bars and grouped into 2D and 3D datasets for each task.
Appendix C Benchmark and score details
C.1 Benchmarks
In this study, we introduce a series of benchmarks designed to evaluate the performance of VLMs on various visual reasoning tasks, including visual search, counting, scene description, and spatial reasoning. Each benchmark was specifically developed to provide rigorous and systematic assessments of different VLM capabilities, utilizing both synthetic and real-world data. These benchmarks are structured to facilitate controlled experiments while maintaining diversity in task difficulty, scene configurations, and object arrangements. Below, we provide detailed descriptions of each benchmark used in our evaluation.
C.1.1 Visual Search Benchmark
@@
C.1.2 Counting Benchmark
To assess the counting ability of VLMs, we developed a comprehensive benchmark based on the generation methodology outlined in [9].
@@ C.1.2 Counting Benchmark
C.1.3 Scene Description Benchmark
@@
To evaluate scene description capabilities in VLMs, we developed a comprehensive benchmark using the feature triplet methodology from [9] with controlled object counts and systematically varied triplet configurations.
@@
This approach enables systematic assessment of scene description performance across: (1) controlled object counts, (2) graduated levels of feature interference through triplet counts, and (3) both 2D and 3D representations. The detailed results for 10, 15, and 20 objects and triplet configurations are presented in Figures  78, and 9, respectively.
@@
C.1.4 Spatial Relationship Benchmark
To evaluate spatial reasoning in VLMs, we developed a comprehensive benchmark using the generation methodology from [9] complemented by natural images.
@@ 3. F1 Score (Reasoning Trace Evaluation)
[…]
@@ 5. Mean Squared Error (Counting)
Our method demonstrates consistent error reduction compared to baseline performance. The most notable improvements occur in natural image scenarios, where exact count prediction remains challenging but our approach achieves substantial MSE reduction (Claude3.5-sonnet: 42.20 versus 1048.63). Similar improvements are evident in synthetic environments, with GPT-4o (1.33 versus 7.50 in 2D) and Qwen2.5-VL (3.89 versus 15.03 in 2D; 6.21 versus 22.04 in 3D) showing significant decreases in squared error.
These MSE results complement our accuracy findings by demonstrating that our method not only increases correct predictions but also reduces the severity of counting errors when they occur. The notable MSE improvement on natural images, despite the inherent difficulty of exact counting in real-world scenes, highlights our method’s ability to prevent large counting deviations.
Table 6: Mean squared error (MSE) of counting task across GPT-4o, Claude3.5-sonnet, LLaMa4, and Qwen2.5-VL on 2D, 3D and natural image datasets using base models and our method.
@@
Simple Ours Simple Ours Simple Ours Simple Ours
@@
Appendix D Implementation Details
In this section, we describe the prompts used across our four evaluation tasks. While we maintain a consistent prompt template for all models, each task requires a tailored instantiation to address its specific objectives. In Figures 10, 11, 12, and 13, we detail the exact prompt formulations employed for each task, which produce the results reported in the Experiments section. Additionally, to assess the model’s Chain-of-Thought (CoT) reasoning ability, we prepend “Let’s think step by step …” to each prompt, encouraging the model to articulate intermediate reasoning before arriving at its final prediction.
@@ Appendix D Implementation Details
Figure 11: Displaying the prompt used for the 2D scenes in the counting task. The phrase “Scan sequentially based on horizontal lines exist in the image” is added to our prompts, in contrast to the simple input.
@@
Appendix E Input and Output examples
In this section, we provide a series of examples to demonstrate the outputs of our row-wise structure across different tasks, including visual search, counting, scene description, and spatial relationship analysis. Each task is illustrated with both synthetic (2D and 3D) and, where applicable, real-world scenes. For each task, we compare the outputs of applying our method with those of a baseline model, highlighting the impact of our approach on a VLM.
E.1 Visual Search
For the visual search task, we present input and output examples from both 2D and 3D datasets (shown in Figures 14 and 15, respectively). In the first example (Figure 14), the objective was to locate the green ’L’ among a set of ’L’ and ’T’ letters within the scene. The baseline model failed to detect the target, returning a [False] output. In contrast, applying row-wise structure successfully identified the green ’L’ and returned a [True] output, demonstrating its effectiveness. This pattern was similarly observed in the 3D scene, where the baseline model again failed to locate the target object, while applying our method accurately detected the presence of the described object within the scene.
E.2 Counting
For the counting task, in addition to the 2D and 3D scene examples, we incorporate a sample from the natural-image dataset utilized in the evaluation of “Learning to Count Everything” [32]. The corresponding examples are presented in Figures 16, 17, and 18, respectively. In all three cases, the VLM model misses some objects or hallucinates extra objects and inaccurately counts the number of objects. In contrast, when our proposed method is applied, the VLM model provides the correct object count.
E.3 Scene Description
In Figures 19 and 20, we present two synthetic examples (one from a 2D scene and the other from a 3D scene) both relevant to the scene description task. The VLM initially fails to accurately identify all objects and their respective attributes within the scenes. However, by applying our proposed method, we improve the model’s performance, enabling it to more accurately and precisely describe the objects in the scenes.
E.4 Spatial Relationship

### v4

@@ 5 Discussion
This research investigated the binding problem in LVLMs, a key factor limiting their performance on structured visual reasoning tasks. We introduced VISER, a method that combines simple visual scaffolding, using horizontal lines, with a targeted textual prompt to encourage sequential and spatially aware processing. Our experiments demonstrate that this external intervention leads to substantial and consistent performance improvements across visual search, counting, scene description, and spatial relationship understanding. Notably, VISER operates in a fully input-agnostic manner, applying the same structure across visual inputs from diverse tasks without additional image processing. These results support the hypothesis that explicit low-level general visual structure improves feature-object binding.
@@
A central finding is the critical role of visual structures. While textual prompts alone proved insufficient and sometimes harmful, the addition of visual lines significantly improved reasoning accuracy. These findings suggest that current LVLMs, despite their advanced linguistic capabilities, can benefit greatly from external visual cues that facilitate a more systematic parsing of complex scenes. The visual structure appears to help models approximate a serial attentional process, mitigating the interference and illusory conjunctions when processing multiple objects and their features in parallel. This localized, sequential processing may also reduce representational interference by effectively managing feature entropy within sub-regions of the image, a factor highlighted in [9] as critical for robust visual reasoning. This supports the idea that addressing the binding problem may require interventions at the level of visual input processing, rather than relying solely on linguistic instruction.
@@
VISER offers a practical and efficient intervention, readily applicable due to its simplicity, negligible computational overhead, and single-query operation, without reliance on external tools, model fine-tuning, or architectural changes (see Appendix G).. This highlights visual input structuring as a key strategy for enhancing LVLM reasoning, beyond linguistic instruction. While effective, the current static nature of the visual scaffolding represents a potential area for refinement (see Appendix E for failure cases). For instance, a fixed scaffolding structure might show reduced gains when the interference between lines and objects is too high, or when object clustering limits the spatial separation benefits of the scaffolds. Future research could investigate adaptive visual scaffolding, where cues are dynamically determined by image content or query specifics. Relatedly, ensemble-based approaches that combine diverse scaffolds (e.g., varying line positions, density, or color) could enhance robustness and reduce sensitivity to edge cases. Moreover, developing LVLM architectures that inherently support serial, spatially grounded attention, perhaps inspired by these findings, could offer more integrated and robust solutions to the binding problem. Extending VISER to other visual tasks, such as hallucination detection and mitigation, also represents a promising direction.
@@
Last but not least, we emphasize that improved visual reasoning in AI systems can benefit a wide range of applications, from assistive technologies and robotics to education and healthcare. However, as LVLMs become more capable, their potential misuse (for instance, in surveillance, automated decision-making, or misinformation) also grows. Our approach does not require model retraining, making it widely accessible, but this also implies that such modifications could be adversarially deployed in black-box systems without transparency or oversight. To mitigate risks, we advocate for responsible use of these techniques in alignment with accountability and transparency guidelines. Future research should profoundly study the unintended biases that scaffolding might introduce in tasks other than spatial reasoning, especially when used in safety-critical applications.
@@ Appendix A Ablation study
A.1 Visual Structures
In our ablation study, we assess multiple visual-structuring strategies and their impact on task performance using the GPT-4o [39] model. All experiments are conducted on a consistent subset of our 2D synthetic dataset [9], with varying object counts per scene. The results in Table 7 show the average performance across these scenes.
Table 7: Comparison of VISER and other ablation methods on GPT-4o in 2D scenes across multiple tasks.
Method Scene Desc. (Edit Dist.) Counting (Accuracy) Visual Search (Accuracy) Spatial Rel. (Accuracy)
Baseline 1.81 16.85 0.48 39.00
Row 1.61 31.55 0.73 44.50
Column 2.00 16.68 0.76 43.50
Grid 2.43 17.40 0.72 51.50
Row–NoRowNum 1.71 19.14 0.70 41.00
The baseline configuration prompts the model to describe the image without any additional structuring cues. As shown, this baseline yields the lowest or near the lowest accuracy in counting and visual search, indicating that unstructured prompts are insufficient for supporting systematic reasoning.
@@ A.1 Visual Structures
In contrast, our row-based horizontal structuring, which explicitly divides the image into horizontal bands, leads to improved performance across most tasks. This design encourages the model to scan the scene in a consistent, line-by-line fashion. Furthermore, removing numerical indices from rows (VISER–NoRowNum) results in a substantial drop in accuracy, highlighting the value of explicit ordering cues in guiding sequential attention.
For the spatial relation task, we chose a grid because the answer space is symmetric across four directions (left, right, above, below), making the grid a more natural fit. However, as shown in Table 7, horizontal lines also perform well in this task. Specifically, for the synthetic spatial relation task with GPT-4o, accuracy improves from 39.00% (baseline) to 44.50% (row method) and 51.50% (grid method). Similarly, in the natural spatial relation task, accuracy increases from 69.39% to 76.92% and 77.43%, respectively. This demonstrates that our simple row-based method is also effective for spatial relations, but the grid method yields the best results.
These results collectively underscore the importance of structured, order-aware visual annotations in enabling LVLMs to reason more effectively over complex spatial arrangements.
A.2 Hyperparameters
In this section, we examined the effect of key scaffold parameters, focusing on the number and thickness of horizontal lines used to segment the scene. The initial configuration applied three horizontal lines of 1 px thickness for all tasks, following the heuristic that too few lines (e.g., 1–2) yield overly coarse partitions with limited benefit, whereas too many lines introduce visual clutter and potential interference artifacts. To systematically assess this, we varied the number of lines (1–12; Figure 5) and the line thickness (1–5 px; Figure 6) across three tasks—Counting, Visual Search, and Scene Description—using two representative models: GPT-4o (closed-source) and Qwen2.5-VL (open-source). Each configuration was evaluated on 300 samples per task.
Across both models, performance dropped at the extremes but remained stable in the 2–6 line range. A similar pattern was observed for line thickness, with consistently strong results up to 5 px. Task-specific effects also emerged: in Scene Description, increasing the number of lines often raised the edit distance (lower is better), indicating that excessive segmentation can hinder coherent free-form output. Overall, these results validate our original heuristic and show that the method is robust to moderate changes in scaffold parameters. While the optimal configuration can vary slightly by task or model, substantial gains are achievable without extensive hyperparameter tuning.
Figure 5: Performance of GPT-4o and Qwen2.5-VL across different tasks (Counting, Visual Search, and Scene Description) with varying numbers of horizontal lines in the input. Accuracy is reported for Counting and Visual Search, while edit distance is used for Scene Description. Baseline represents performance with no horizontal lines.
Figure 6: Performance of GPT-4o and Qwen2.5-VL across different tasks with varying line thicknesses (in pixels) introduced by VISER. Each column represents results for a specific line width, ranging from 1 px to 5 px. Accuracy is reported for the Counting and Visual Search tasks, while lower edit distance indicates better performance for the Scene Description task. The Baseline corresponds to performance without added lines (i.e., line thickness of zero).
A.3 Imaginary Lines
To evaluate the impact of prompting the model to imagine auxiliary structures (e.g., horizontal lines) without rendering them, we introduced an imaginary-baseline variant. This baseline was prompted to scan the image row by row using imaginary lines—similar to the setup illustrated in Figure 27—but without explicitly drawing them. As shown in Table 8, VISER outperforms the imaginary-baseline across all tasks: achieving an 8.5% gain in spatial relation accuracy, a 13.31% improvement in counting, a 0.12 reduction in edit distance for scene description, and a 5.00% increase in harmonic mean accuracy for visual search.
Table 8: Performance of the baseline when prompted to imagine horizontal lines (as in Figure 27), evaluated across tasks using the GPT-4o model.
Method Visual Search (Acc.) Counting (Acc.) Scene Desc. (Edit Dist.) Spatial Rel. (Acc.)
Baseline 48.00 16.85 1.84 43.00
Imaginary-Baseline 68.00 18.24 1.75 44.00
VISER 73.00 31.55 1.63 52.50
A.4 Reasoning trace evaluation
@@
Figure 7: Comparison between VISER and GPT-4o in the 2D scene description task, where models sequentially scan the image row by row to identify objects. Results are reported for scenes with 10, 15, and 20 objects using three evaluation metrics: F1 score, Jaccard similarity, and Edit distance.
Appendix B Benchmark and score details
B.1 Benchmarks
In this study, we introduce a series of benchmarks designed to evaluate the performance of LVLMs on various visual reasoning tasks, including visual search, counting, scene description, and spatial reasoning. Each benchmark was specifically developed to provide rigorous and systematic assessments of different VLM capabilities, utilizing both synthetic and real-world data. These benchmarks are structured to facilitate controlled experiments while maintaining diversity in task difficulty, scene configurations, and object arrangements. Below, we provide detailed descriptions of each benchmark used in our evaluation.
B.1.1 Visual Search Benchmark
@@
B.1.2 Counting Benchmark
To assess the counting ability of LVLMs, we developed a comprehensive benchmark based on the generation methodology outlined in [9].
@@ B.1.2 Counting Benchmark
B.1.3 Scene Description Benchmark
@@
To evaluate scene description capabilities in LVLMs, we developed a comprehensive benchmark using the feature triplet methodology from [9] with controlled object counts and systematically varied triplet configurations.
@@
This approach enables systematic assessment of scene description performance across: (1) controlled object counts, (2) graduated levels of feature interference through triplet counts, and (3) both 2D and 3D representations. The detailed results for 10, 15, and 20 objects and triplet configurations are presented in Figures  89, and 10, respectively.
@@
B.1.4 Spatial Relationship Benchmark
To evaluate spatial reasoning in LVLMs, we developed a comprehensive benchmark using the generation methodology from [9] complemented by natural images.
@@ 3. F1 Score (Reasoning Trace Evaluation)
$\text{F1}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (4)
@@ 5. Mean Squared Error (Counting)
VISER demonstrates consistent error reduction compared to baseline performance. The most notable improvements occur in natural image scenarios, where exact count prediction remains challenging but our approach achieves substantial MSE reduction (Claude3.5-sonnet: 42.20 versus 1048.63). Similar improvements are evident in synthetic environments, with GPT-4o (1.33 versus 7.50 in 2D) and Qwen2.5-VL (3.89 versus 15.03 in 2D; 6.21 versus 22.04 in 3D) showing significant decreases in squared error.
These MSE results complement our accuracy findings by demonstrating that VISER not only increases correct predictions but also reduces the severity of counting errors when they occur. The notable MSE improvement on natural images, despite the inherent difficulty of exact counting in real-world scenes, highlights VISER’s ability to prevent large counting deviations.
Table 9: Mean squared error (MSE) of counting task across GPT-4o, Claude3.5-sonnet, LLaMa4, and Qwen2.5-VL on 2D, 3D and natural image datasets using base models and VISER.
@@
Baseline VISER Baseline VISER Baseline VISER Baseline VISER
@@
B.3 Performance Variance and Statistical Significance
Each reported score in our experiments is averaged over a substantial number of samples (typically 50–100 per configuration), which provides a stable estimate of performance. To minimize non-determinism, all evaluations use greedy decoding (temperature = 0), ensuring deterministic outputs for each input. This reduces variance introduced by stochastic generation and makes the results reproducible. While we do not repeat each setup multiple times, we emphasize robustness by evaluating across a broad set of conditions, including four tasks, both 2D and 3D synthetic data, real-world datasets, and multiple open- and closed-source models.
Across 120 pairwise comparisons between VISER and the baseline (excluding supplementary material), VISER outperforms in 103 cases, shows decreases in 17 cases, and yields ties in 7 cases. Most drops occur with LLaMA4, a model that generally performs poorly on these tasks. Considering only non-tied cases (96 wins vs. 17 losses), a one-sided binomial sign test yields a p-value of $7.4\times 10^{-15}$, strongly rejecting the null hypothesis that VISER is no better than the baseline. The tied cases typically correspond to trivial scenarios with zero or saturated performance.
Appendix C Implementation Details
In this section, we describe the prompts used across our four evaluation tasks. While we maintain a consistent prompt template for all models, each task requires a tailored instantiation to address its specific objectives. In Figures 11, 12, 13, and 14, we detail the exact prompt formulations employed for each task, which produce the results reported in the Experiments section. Additionally, to assess the model’s Chain-of-Thought (CoT) reasoning ability, we prepend “Let’s think step by step …” to each prompt, encouraging the model to articulate intermediate reasoning before arriving at its final prediction.
@@ Appendix C Implementation Details
Figure 12: Displaying the prompt used for the 2D scenes in the counting task. The phrase “Scan sequentially based on horizontal lines exist in the image” is added to our prompts, in contrast to the baseline input.
@@
Appendix D Input and Output examples
In this section, we provide a series of examples to demonstrate the outputs of our row-wise structure across different tasks, including visual search, counting, scene description, and spatial relationship analysis. Each task is illustrated with both synthetic (2D and 3D) and, where applicable, real-world scenes. For each task, we compare the outputs of applying VISER with those of a baseline model, highlighting the impact of our approach on a VLM.
D.1 Visual Search
For the visual search task, we present input and output examples from both 2D and 3D datasets (shown in Figures 15 and 16, respectively). In the first example (Figure 15), the objective was to locate the green ’L’ among a set of ’L’ and ’T’ letters within the scene. The baseline model failed to detect the target, returning a [False] output. In contrast, applying row-wise structure successfully identified the green ’L’ and returned a [True] output, demonstrating its effectiveness. This pattern was similarly observed in the 3D scene, where the baseline model again failed to locate the target object, while applying VISER accurately detected the presence of the described object within the scene.
D.2 Counting
For the counting task, in addition to the 2D and 3D scene examples, we incorporate a sample from the natural-image dataset utilized in the evaluation of “Learning to Count Everything” [32]. The corresponding examples are presented in Figures 17, 18, and 19, respectively. In all three cases, the VLM model misses some objects or hallucinates extra objects and inaccurately counts the number of objects. In contrast, when our proposed method is applied, the VLM model provides the correct object count.
D.3 Scene Description
In Figures 20 and 21, we present two synthetic examples (one from a 2D scene and the other from a 3D scene) both relevant to the scene description task. The VLM initially fails to accurately identify all objects and their respective attributes within the scenes. However, by applying our proposed method, we improve the model’s performance, enabling it to more accurately and precisely describe the objects in the scenes.
D.4 Spatial Relationship